# A Test of Time: Predicting the Sustainable Success of Online Collaboration in Wikipedia

### v1

@@ 5. Experimental Setup
5.1. Classification Models
We experiment with an array of algorithms ranging from simple logistic regression to more sophisticated neural networks. We do not expect complex deep learning and transformer architectures to outperform standard ML models in our settings, since we formulate SustainPedia as tabular data (Shwartz-Ziv and Armon, 2022). [^7] We do take linguistic and time-related features into account, but formulate those as tabular features.
We report only the results obtained from the Gradient-Boosted-Decision-Trees (GBT) model (Friedman, 2002), which achieved comparable results to other ensemble models such as Random Forest (Breiman, 2001) and CatBoost (Dorogush et al., 2018). We use a deviance loss function for the GBT model. We use the sklearn (Pedregosa et al., 2011) implementation of the algorithm, with the default parameters of the package. Specifically, we build the model with 100 estimators (i.e., trees). We also tested different approaches to optimizing some of the model parameters.
@@ 5.2. Evaluation
6. Results and Analysis
Table 2. Modeling results. A compositional model of all features outperforms other alternatives in both $FA$ and $GA$ use cases. We report the average measures (±$\pm std$) over 100 resampled bootstrap iterations. We calculate the measures over the positive class (unsustainable articles) besides Macro-F1. The baseline is a model that uses a single feature – Num-of-Revisions-Normalized.
@@ 6. Results and Analysis
Using a single feature type, the highest Macro F1 and AU-ROC are obtained by the User Experience feature set. This holds for both $FA$ and $GA$ use cases. This highlights the importance of intellectual capital while building a collaborative team (Nahapiet and Ghoshal, 1998; Coff, 1997). Specifically, having Wikipedians who are experienced in editing and contributing to successful projects is crucial in creating sustainable articles. We note that while trailing behind, the Edit History feature set alone achieved a decent ($\thicksim 15\%$) improvement over the baseline model. On the other hand, the Topics feature set, although comprised of a large set of features, does not outperform other models.
@@
Figure 2. Models performance. We present the aggregate values over 100 resampled bootstrap iterations. Boxplots represent the interquartile range (IQR). Horizontal dashed lines indicate the median values. The black upward (downward) triangle marks the average value of the $FA$ ($GA$) use case. AU-ROC y-axis is the Area under the Receiver-Operator. A random model achieves a 0.5 AU-ROC.
6.1. Association Heatmaps
Figure 3. Heat maps of feature interactions with unsustainability. Unsustainability is distributed among feature pairs in insightful ways. The value in each cell is the mean unsustainability per bin. We omit presenting values where the number of cases is too small (¡10).
@@ 6.1. Association Heatmaps
The trade-off between the number of contributors and how work is distributed between them has been long questioned (Ortega et al., 2008; Arazy and Nov, 2010). Figure 3a highlights this balance for our study. It is clear that a lower Gini (unequal contribution) among editors is strongly associated with unsustainability. We also observe that having (too) many collaborators is associated with unsustainability.
@@
Figure 3c highlights that many editors expressing negative sentiment in talk pages put the article at risk of unsustainable status. Few editors expressing positive sentiment in talk pages is more likely to produce sustainable articles. This expands the existing understanding of the relation between discussion patterns and quality assessment in Wikipedia (Kittur and Kraut, 2008; Stvilia et al., 2005).
6.2. Error Analysis
@@
Figure 4. Average number of reviews per population, $FA$ use case. The FP population is reviewed significantly more frequently compared to the sustainable articles population, which highlights their demotion potential – identified by the Wikipedia community. Naturally, the number of reviews for the unsustainable population is the highest since these demoted articles go over at least a single review.
@@ 6.3. Corpus Size
6.4. Model Interpretation
To support RQ2, we use the SHAP framework (Lundberg and Lee, 2017) to shed light on the specific features that impact the model’s prediction. A SHAP value is calculated for each explanatory variable (feature) and input instance (i.e., an article in our case) for a specific model. A high (low) SHAP value indicates the positive (negative) impact of the feature on the prediction of the specific instance. Looking at the aggregate values for a specific feature provides a way to understand feature importance and their positive/negative impact.
@@ 6.4. Model Interpretation
Each of the three last features comes from a different feature set. Median-Num-of-Discussers highlights that (too) many Wikipedians participating in the talk page discussions increase the likelihood of an article being unsustainable. This observation is consistent with previous research on crowd collaboration success (Brooks Jr, 1995; Kittur and Kraut, 2008), which emphasizes that an increase in the number of contributors does not inherently improve the quality of the collaboration but rather requires different coordination strategies. We assume that the high association of the Politics topic with unsustainability is due to (i) Many controversial issues raised about political items (Rad and Barbosa, 2012), and (ii) The frequent information updates and new events in the political domain, leading articles becoming quickly outdated and thus being more difficult to sustain.
Figure 5. Top ten SHAP features importance. Almost all feature sets are represented in the top ten, while the Experience set dominates the list. Orange (blue) bars correspond to negative (positive) impact on the prediction of unsustainability.
7. Discussion and Limitations
@@
Quantifying the sustainable success of crowd collaborations goes beyond Wikipedia articles and extends to other online and offline collaborative efforts such as crowdsourced open-source software and entrepreneurial teams (Lechler, 2001; Lazar et al., 2020; Daniel et al., 2018; Saadat et al., 2020). In such collaborations, success is not merely determined by the short-term success or the initial hype they make but rather by the longevity and lasting impact they generate. The social dynamics driving these collaborations share parallels with the challenges faced in maintaining the long-term quality and relevance of Wikipedia articles. We see the importance of our study also in introducing a framework to measure and model the sustainable success of online/offline human crowd collaborations beyond Wikipedia.
@@
This study focuses exclusively on English Wikipedia articles, limiting its scope to non-English projects. Additionally, the data used is restricted to the point in time when the study was conducted, meaning that articles may be demoted in the future, which we cannot account for. Lastly, our analysis is based solely on Wikipedia’s internal dynamics, without incorporating interactions that may occur outside of the platform (e.g., social media).
8. Summary
@@ Appendix A Appendix
Figure 6. Review process alert of the Durian article. The screenshot was taken from an April 2022 version of the talk page. Unfortunately, the article was demoted in June 2022 after this review process. We conceal all identities to preserve editors’ right to remain anonymous.
Figure 7. Average number of reviews per population, $FA$ use case. The FP population includes only the top-100 articles predicted by the model to be unsustainable (i.e., articles are sorted by the predicted likelihood of the model). The FP population is reviewed significantly more frequently compared to the sustainable articles population, which highlights their demotion potential identified by the Wikipedia community. This is aligned with the pattern we illustrate in Figure 4 when taking the whole FP population (and not only the top-100 articles). Naturally, the number of reviews for the unsustainable population is the highest since these articles go over at least a single review before being demoted.
Table 3. Discussion features related to the structure of discussions. Details about this feature set are provided in Section 4.3.
@@
Table 4. Population-level results based bootstrapped scores across 100 model runs. We report the average measures (±$\pm std$) over 100 resampled bootstrap iterations. We calculate all measures over the positive class (unsustainable articles) besides Macro-F1 and AU-ROC. ‘AU-ROC’ is the Area under the Receiver-Operator Curve. ‘Prec@X’ is the Precision at X percentage of the most likely unsustainable articles by the model. The baseline model uses a single explanatory feature – Num-of-Revisions-Normalized. $\uparrow$ Indicates that a higher value is better.
@@
Table 5. Five-Fold-CV modeling results. We report the average measures (±$\pm std$) over five-stratified-fold-CVs. We calculate all measures over the minority class (unsustainable articles) besides Macro-F1 and AU-ROC. ‘AU-ROC’ is the Area under the Receiver-Operator Curve. ‘Prec@X’ is the Precision at X percentage of the most likely unsustainable articles by the model. The baseline model uses a single explanatory feature – Num-of-Revisions-Normalized. $\uparrow$ Indicates that a higher value is better.
@@
Table 6. The top ten False Positive (FP) articles. These articles have the highest unsustainability (demotion) potential according to our model (i.e., articles are sorted by the predicted likelihood of the model). ‘Pred.’ represents the model’s predicted likelihood that the article will become unsustainable.
@@
Table 7. Articles that have been recently promoted to an $FA$ status (01/2019-09/2023), which have the highest potential to be unsustainable according to our model (i.e., articles are sorted by the predicted likelihood of the model). ‘Pred.’ represents the model’s predicted likelihood that the article will become unsustainable.
@@
Figure 8. The model’s performance improves as the corpus size gets larger and the model is exposed to more positive (unsustainable) articles. Performance improvement continues to increase and has not saturated. Improvement is measured as a percentage compared to a random model. As the corpus expands, the model shows better results in both measures (right and left y-axis). The dashed line corresponds to values on the right y-axis. While error bars are included, some are too small to be visually noticeable.

### v2

@@ 5 Experimental Setup
5.1 Classification Models
We experiment with an array of algorithms ranging from simple logistic regression to more sophisticated neural networks. We do not expect complex deep learning and transformer architectures to outperform standard ML models in our settings, since we formulate SustainPedia as tabular data Shwartz-Ziv and Armon (2022). [^7] We do take linguistic and time-related features into account, but formulate those as tabular features.
We report only the results obtained from the Gradient-Boosted-Decision-Trees (GBT) model Friedman (2002), which achieved comparable results to other ensemble models such as Random Forest Breiman (2001) and CatBoost Dorogush et al. (2018). We use a deviance loss function for the GBT model. We use the sklearn Pedregosa et al. (2011) implementation of the algorithm, with the default parameters of the package. Specifically, we build the model with 100 estimators (i.e., trees). We also tested different approaches to optimizing some of the model parameters.
@@ 5.2 Evaluation
6 Results and Analysis
Table 2: Modeling results. A compositional model of all features outperforms other alternatives in both $FA$ and $GA$ use cases. We report the average measures (±$\pm std$) over 100 resampled bootstrap iterations. We calculate the measures over the positive class (unsustainable articles) besides Macro-F1. The baseline is a model that uses a single feature – Num-of-Revisions-Normalized.
@@ 6 Results and Analysis
Using a single feature type, the highest Macro F1 and AU-ROC are obtained by the User Experience feature set. This holds for both $FA$ and $GA$ use cases. This highlights the importance of intellectual capital while building a collaborative team Nahapiet and Ghoshal (1998); Coff (1997). Specifically, having Wikipedians who are experienced in editing and contributing to successful projects is crucial in creating sustainable articles. We note that while trailing behind, the Edit History feature set alone achieved a decent ($\thicksim 15\%$) improvement over the baseline model. On the other hand, the Topics feature set, although comprised of a large set of features, does not outperform other models.
@@
Figure 2: Models performance. We present the aggregate values over 100 resampled bootstrap iterations. Boxplots represent the interquartile range (IQR). Horizontal dashed lines indicate the median values. The black upward (downward) triangle marks the average value of the $FA$ ($GA$) use case. AU-ROC y-axis is the Area under the Receiver-Operator. A random model achieves a 0.5 AU-ROC.
6.1 Association Heatmaps
Figure 3: Heat maps of feature interactions with unsustainability. Unsustainability is distributed among feature pairs in insightful ways. The value in each cell is the mean unsustainability per bin. We omit presenting values where the number of cases is too small (<10).
@@ 6.1 Association Heatmaps
The trade-off between the number of contributors and how work is distributed between them has been long questioned Ortega et al. (2008); Arazy and Nov (2010). Figure 3a highlights this balance for our study. It is clear that a lower Gini (unequal contribution) among editors is strongly associated with unsustainability. We also observe that having (too) many collaborators is associated with unsustainability.
@@
Figure 3c highlights that many editors expressing negative sentiment in talk pages put the article at risk of unsustainable status. Few editors expressing positive sentiment in talk pages is more likely to produce sustainable articles. This expands the existing understanding of the relation between discussion patterns and quality assessment in Wikipedia Kittur and Kraut (2008); Stvilia et al. (2005).
6.2 Error Analysis
@@
Figure 4: Average number of reviews per population, $FA$ use case. The FP population is reviewed significantly more frequently compared to the sustainable articles population, which highlights their demotion potential – identified by the Wikipedia community. Naturally, the number of reviews for the unsustainable population is the highest since these demoted articles go over at least a single review.
@@ 6.3 Corpus Size
6.4 Model Interpretation
To support RQ2, we use the SHAP framework Lundberg and Lee (2017) to shed light on the specific features that impact the model’s prediction. A SHAP value is calculated for each explanatory variable (feature) and input instance (i.e., an article in our case) for a specific model. A high (low) SHAP value indicates the positive (negative) impact of the feature on the prediction of the specific instance. Looking at the aggregate values for a specific feature provides a way to understand feature importance and their positive/negative impact.
@@ 6.4 Model Interpretation
Each of the three last features comes from a different feature set. Median-Num-of-Discussers highlights that (too) many Wikipedians participating in the talk page discussions increase the likelihood of an article being unsustainable. This observation is consistent with previous research on crowd collaboration success Brooks Jr (1995); Kittur and Kraut (2008), which emphasizes that an increase in the number of contributors does not inherently improve the quality of the collaboration but rather requires different coordination strategies. We assume that the high association of the Politics topic with unsustainability is due to (i) Many controversial issues raised about political items Rad and Barbosa (2012), and (ii) The frequent information updates and new events in the political domain, leading articles becoming quickly outdated and thus being more difficult to sustain.
While not all the dominant features we identified are actionable, many are. Notably, a gradual progression toward attaining recognized high-quality status was found to be strongly associated with sustainable success. This finding suggests implementing a policy of incremental quality assessments, rather than immediate promotion to the highest assessment level, may be beneficial. In the context of Wikipedia, this approach is exemplified by initially designating articles as $GA$ before considering them for an $FA$ status. Additionally, our analysis highlights the critical role of experienced collaborators with a demonstrated track record of success. This actionable feature should motivate collaborative teams to engage with experienced users and convince them to join their collaborative effort to enhance the likelihood of achieving sustainable success.
Figure 5: Top ten SHAP features importance. Almost all feature sets are represented in the top ten, while the Experience set dominates the list. Orange (blue) bars correspond to negative (positive) impact on the prediction of unsustainability.
7 Discussion and Limitations
Other studies of collaborative success
The objective of most existing studies is ‘Success Prediction.’ That is, to research what leads crowd collaboration to initial success. Our objective is different. We aim to measure these collaborations through the lens of whether success holds over time, using our new ‘Sustainable Success’ measure. We see these two objectives as different by nature. However, highlighting the similar (and different) dominant features between ‘Success Prediction’ and ‘Sustainability Prediction’ is insightful.
A key feature that we find to be associated with ‘Sustainability Prediction’ is a prolonged and gradual path toward attaining a recognized high-quality status. This feature is unique to our study and has yet to be recognized in other ‘Success Prediction’ studies. [^11] Naturally, ‘Success Prediction’ studies are limited in the way this feature can be constructed as their target feature is different than ours. The participation of experienced team members with a proven history of involvement in successful collaborations, which we find a significant feature, was also found to be significant in a few ‘Success Prediction’ studies Stein and Hess (2007); Betancourt et al. (2016).
For other features, such as number-of-editors, editors-gini, and most network features, while we did not find those to be the most dominant, other ‘Success Prediction’ studies found them as significant predictors Arazy and Nov (2010); Qin and Cunningham (2012); Raman et al. (2020); Lin and Wang (2020).
Overall, our findings emphasize the unique features critical to sustainable success, while also identifying shared elements with existing research on the general task of success prediction.
@@ 7 Discussion and Limitations
We use Wikipedia as a testbed to answer our research questions. However, quantifying the sustainable success of crowd collaborations goes beyond Wikipedia articles and extends to other online and offline collaborative efforts such as crowdsourced open-source software and entrepreneurial teams Lechler (2001); Lazar et al. (2020); Daniel et al. (2018); Saadat et al. (2020). These crowd-collaboration projects span multiple domains and play a critical role in large-scale projects.
In such collaborations, success is not merely determined by the short-term success or the initial hype they make but rather by the longevity and lasting impact they generate. The social dynamics driving these collaborations share parallels with the challenges faced in maintaining the long-term quality and relevance of Wikipedia articles. We see the importance of our study also in introducing a framework to measure and model the sustainable success of online/offline human crowd collaborations beyond Wikipedia.
@@
This study focuses exclusively on English Wikipedia articles, limiting its scope to non-English projects. Additionally, the data used is restricted to the point in time when the study was conducted, meaning that articles may be demoted in the future, which we cannot account for. As described in Section 4, our analysis is based solely on Wikipedia’s internal dynamics, without incorporating interactions that may occur outside of the platform (e.g., social media).
Finally, our research is based on “naturally occurring” data and events, rather than a “designed experiment”. Since we do not control the available data or its scale, this imposes limits on our ability to make causal claims—particularly in drawing causal conclusions between explanatory factors and sustainable success. Future research that incorporates controlled experiments could provide deeper insights into the causal relationships between key factors and sustainable success, allowing for a more rigorous evaluation of their long-term impact.
8 Summary
@@ Appendix A Appendix
Figure 6: Review process alert of the Durian article. The screenshot was taken from an April 2022 version of the talk page. Unfortunately, the article was demoted in June 2022 after this review process. We conceal all identities to preserve editors’ right to remain anonymous.
Figure 7: Average number of reviews per population, $FA$ use case. The FP population includes only the top-100 articles predicted by the model to be unsustainable (i.e., articles are sorted by the predicted likelihood of the model). The FP population is reviewed significantly more frequently compared to the sustainable articles population, which highlights their demotion potential identified by the Wikipedia community. This is aligned with the pattern we illustrate in Figure 4 when taking the whole FP population (and not only the top-100 articles). Naturally, the number of reviews for the unsustainable population is the highest since these articles go over at least a single review before being demoted.
Table 3: Discussion features related to the structure of discussions. Details about this feature set are provided in Section 4.3.
@@
Table 4: Population-level results based bootstrapped scores across 100 model runs. We report the average measures (±$\pm std$) over 100 resampled bootstrap iterations. We calculate all measures over the positive class (unsustainable articles) besides Macro-F1 and AU-ROC. ‘AU-ROC’ is the Area under the Receiver-Operator Curve. ‘Prec@X’ is the Precision at X percentage of the most likely unsustainable articles by the model. The baseline model uses a single explanatory feature – Num-of-Revisions-Normalized. $\uparrow$ Indicates that a higher value is better.
@@
Table 5: Five-Fold-CV modeling results. We report the average measures (±$\pm std$) over five-stratified-fold-CVs. We calculate all measures over the minority class (unsustainable articles) besides Macro-F1 and AU-ROC. ‘AU-ROC’ is the Area under the Receiver-Operator Curve. ‘Prec@X’ is the Precision at X percentage of the most likely unsustainable articles by the model. The baseline model uses a single explanatory feature – Num-of-Revisions-Normalized. $\uparrow$ Indicates that a higher value is better.
@@
Table 6: The top ten False Positive (FP) articles. These articles have the highest unsustainability (demotion) potential according to our model (i.e., articles are sorted by the predicted likelihood of the model). ‘Pred.’ represents the model’s predicted likelihood that the article will become unsustainable.
@@
Table 7: Articles that have been recently promoted to an $FA$ status (01/2019-09/2023), which have the highest potential to be unsustainable according to our model (i.e., articles are sorted by the predicted likelihood of the model). ‘Pred.’ represents the model’s predicted likelihood that the article will become unsustainable.
@@
Figure 8: The model’s performance improves as the corpus size gets larger and the model is exposed to more positive (unsustainable) articles. Performance improvement continues to increase and has not saturated. Improvement is measured as a percentage compared to a random model. As the corpus expands, the model shows better results in both measures (right and left y-axis). The dashed line corresponds to values on the right y-axis. While error bars are included, some are too small to be visually noticeable.